1.find the number of a quanta of radiation of frequency 4.67*1013s-1 that must be absorabed in order to melt 5gm of ice. energy required to melt 1gm of ice is 333j?
2.calculate the velocity of electrostatic potential energy of to electrons separated by 3.0 in vacuum. express your answers in jouls and electron volt?
3.energy required to stop the ejection of electrons is 0.24 eV. calculate the work function when radiationsof ?=253.7 nm strikes the plate?

Grade:

1 Answers

Pankaj
9 years ago
Using
We can calculate the energy of one photon using
E = hv
Given that
v = 4.67×1013s-1
h = 6.62 ×10−34 J.S
We get, E = (6.62 ×10−34)( 4.67×1013s-1 ) = 30.94 × 10-21 J
Number of quanta required to melt five gram of ice = 1665/30.94 × 10-21 =
53.81 × 1021